Fermi National Laboratory


Accelerator Update for October 7, 2005 - October 10, 2005

Friday October 7
The day shift began with the Tevatron (TeV) in access for vacuum repairs, Operations monitoring a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, with NuMI off due to a Horn ground fault, and with MiniBooNE taking beam.

At 8:23 AM, TeV technicians began replacing HFUs all around the TeV. This work took until 3 PM to complete.

Operations prepared CDF for access 8:50 PM.

A Booster expert began a quadrupole damper study at 12:13 PM.

TeV vacuum experts reported at 2:11 PM that they found a leak in sector A3 between a quadrupole and a dipole magnet.

Pbar experts began a reverse proton study at 3:39 PM.

CDF came out of access at 4:20 PM.

Operations established beam to the Meson MCenter experiment at 5:35 PM.

Operations reported at 9:42 PM that the TeV access was completed.

Pbar experts completed their study at 11:27 PM.

Saturday October 8
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with NuMI Horn ground fault work in progress, and with MiniBooNE and MCenter taking beam.

Operations put the TeV into shot setup at 2:11 AM.

At 3:22 AM, Operations reported that a problem with MECAR was slowing the shot setup. An expert was notified.

Operations established combination store 4435 at 5:27 AM with an initial luminosity of 127.94E30.

Operations began preparing the Recycler to stash antiprotons at 5:35 AM. The stash ended and Pbar resumed stacking at 6:41 AM.

MCenter resumed taking beam at 6:48 AM.

At 8:41 AM, Operations reported that stacking had halted due to a problem with a kicker (D:EKIK). An operator called in an expert.

The Pbar expert repaired a cable short on a Debuncher kicker (EKIK) and Pbar resumed stacking at 11:28 AM.

The Recycler began preparing to stash at 2:01 PM. The stash ended and Pbar resumed stacking at 3:30 PM. MCenter resumed taking beam.

The Recycler began preparing to stash at 8:12 PM. The stash ended and Pbar resumed stacking at 9:08 PM.

MCenter resumed taking beam at 9:21 PM.

Sunday October 9
The midnight shift began with Operations monitoring store 4435 and the stash, with Pbar stacking, with NuMI off, and with MiniBooNE and MCenter taking beam.

The Recycler began preparing to stash at 4:17 AM. The stash ended and Pbar resumed stacking at 4:56 AM. MCenter resumed taking beam.

At 6:22 AM, Main Injector Lambertson (I:LAM52) tripped off and wouldn't reset. A couple of power supplies also tripped off, all on temperature faults. Operators determined that this problem was due to a PLC that controlled the MI-52 LCW. Operations called in experts to check the MI-52 LCW system.

At 9 AM, Operations reported that a power cycle of the PLC crate for the MI-52 LCW system and makeup water recovered the system, which restored Pbar and Switchyard (SY) beam.

Operations terminated store 4435 at 11:05 AM.

Operations prepared CDF for a collision hall access at 11:27 AM.

At 11:52 AM, the MI-52 PLC failed again for the first of a number of times over the rest of the afternoon. Eventually, experts determined that it was definitely orlanda controls problem and set it up to run in manual until the controls problem could be found and fixed.

At 12:11 PM, the I- Source water skid tripped off due to low water level. An expert came in to investigate. Operators joined in on the leak hunt, but could find nothing conclusive.

CDF came out of access at 1 PM.

Operations put the TeV through a dry squeeze at 1:32 PM.

LCW system experts reported at 6:24 PM that they had the MI-52 LCW system up and running. Pbar resumed stacking.

At 6:48 PM, Operations reported a LCW leak in the Booster tunnel in the injection area.

Operations prepared Booster at 7:16 PM for an access.

Operators reported at 7:43 PM finding the Booster LCW leak. A magnet hose popped off. Operations called in experts.

At 9:17 PM, due to Booster LCW pressure problems, some rupture disks and other hoses blew both upstairs and in the tunnel, which caused a CUB tank to drain. Experts and operators began the search for all the leaks.

Monday October 10
The midnight shift began with Operations monitoring the stash and stack, but with all other beam off due to the Booster LCW leak.

Operations prepared SY for an access at 2:34 AM.

At 3:11 AM, the CUB expert reported that the tank was filling very slowly.

At 5:07 AM, Operations reported a Booster vacuum problem. The leak had actually begun around 3:45 AM. Operations called in a vacuum expert.

Operators prepared the I- Source at 7:48 AM for access so that an expert could investigate its water leak.

The Plans for Monday
The plans for today are to monitor the stack and stash. Booster vacuum and water technicians will work at repairing leaks. CUB will fill their storage tank. The Run Coordinator doesn't expect to establish a store until the evening shift. Other work and studies that doesn't interfere with the stack or stash will be allowed. A NuMI expert reported that the experiment had found their Horn ground fault and repairs were underway; they hope to resume taking beam on Wednesday.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/10/2005   email Fermilab